OpenStudy (anonymous):

If you know that the period of a pendulum is 1.87 seconds, what is the length of that pendulum? (Assume that we are on Earth and that gravity is 9.81 meters/second².) Select one of the options below as your answer: A. 0.87 centimeters B. 2.1 meters C. 1.6 meters D. 0.87 meters E. 8.3 meters

OpenStudy (anonymous):

\[\color{green}{T = 2 \pi \sqrt{\frac{l}{g}}}\] Solve for l.